# TenonOS: A Self-Generating Intelligent Embedded Operating System Framework for Edge Computing

### v1

@@ 4 Related Work
Operating systems designed for edge and embedded environments must carefully balance flexibility, performance, and strong isolation. Meeting these demands necessitates advancements in both operating system architecture and the underlying virtualization framework. Lightweight, application-specific operating systems—such as Library OSes (LibOS)—minimize overhead and enable precise control over resource usage. Concurrently, minimal type-1 hypervisors incorporating novel architectural designs provide robust isolation and predictable performance, both of which are critical for safety-critical and mixed-criticality applications.
@@ 4.1 Type-1 Hypervisors
Virtualization technology has advanced considerably, with type-1 hypervisors playing a key role in balancing performance, security, and resource isolation. Xen [39], a pioneer in type-1 hypervisor design, introduced the classic split-domain architecture. Running directly on bare-metal hardware, Xen separates device management (handled by Dom0) from guest virtual machines (hosted in DomU), ensuring strong isolation and security. Its support for both ARM and x86 architectures enhances compatibility in heterogeneous edge deployments. Extensions such as RT-Xen further enable latency-sensitive applications in industrial automation and IoT environments. Rust-Shyper [40] represents a recent innovation at the programming language level. By leveraging Rust’s memory safety guarantees, Rust-Shyper minimizes vulnerabilities that are common in C/C++ implementations. This approach enhances security and reliability, which is especially important for edge scenarios such as 5G base stations and distributed AI inference.
Further advancing hypervisor architecture, NOVA [38] adopts a microhypervisor (microkernel-inspired) strategy, drastically reducing the trusted computing base (TCB) and improving system modularity. By delegating most functionalities to unprivileged user-space components, NOVA achieves strong fault tolerance, making it suitable for edge systems deployed in untrusted environments like smart city infrastructure and autonomous devices. Continuing this trend towards modularity and determinism, BAO [31] is a lightweight, bare-metal hypervisor designed for multi-core embedded systems. BAO applies a modular architecture and statically allocates hardware resources during initialization, achieving deterministic behavior with strong spatial and temporal isolation. This makes BAO particularly suitable for safety-critical and mixed-criticality applications, such as autonomous vehicles and medical devices.
This progression—from Xen’s classic architecture, through Rust-based language innovation, microkernel modularization in NOVA, to BAO’s lightweight and modular design—demonstrates how virtualization technology is evolving towards greater modularity, security, and efficiency for modern edge computing demands.
Figure 1: The overall architecture of TenonOS. The left section illustrates a traditional monolithic OS-Hypervisor design. The middle section depicts the Librarization process, where system functionalities are modularized into a library pool comprising reusable components. Then dynamic orchestration and composition mechanism allows rebuilt and generate TenonOS to adapt to heterogeneous hardware resources, forming a lightweight and agile LibOS on LibOS structure.
@@ 4.2 Flexibility and Performance of Embedded LibOS
LibOS have emerged as a promising paradigm for constructing flexible, lightweight and high-performance software systems. Unlike traditional monolithic operating systems, a LibOS encapsulates only the essential OS abstractions and services required by an application, enabling application-specific optimizations. Early projects such as Exokernel [41] pioneered the separation of application logic from resource management, allowing applications to directly manage hardware resources.
Recent advancements have focused on integrating LibOS into virtualized environments. For example, Graphene [42] demonstrated the feasibility of running unmodified applications in isolated LibOS instances. This approach maintains compatibility with existing APIs while enhancing security and performance. In cloud computing, LibOS-based systems such as OSv [43] and Unikraft [44] have achieved significant improvements in startup latency, memory footprint, and runtime efficiency compared to traditional virtual machines. By leveraging their isolated architecture, these systems minimize overhead and enable further application-specific optimizations.
@@ 5.1 TenonOS Architecture and Design Philosophy
Figure 1 presents the overall architecture of TenonOS, which treats the virtualization layer and guest OS as a unified whole. Traditionally, as shown on the left of the figure, the hypervisor and operating system are organized in a monolithic, tightly coupled stack. In contrast, TenonOS proposes a novel Librarization process, illustrated in the middle of the figure, where both the hypervisor and the OS are dismantled into a unified pool of reusable library components. These modules encapsulate core functionalities, and a dynamic library orchestration mechanism selects and composes the appropriate modules in response to specific hardware and application contexts. The right side of the figure shows the reconstructed modular stack: the bottom layer, Mortise, is a LibOS-based hypervisor that runs directly on hardware and can act as either a traditional OS or dynamically extend to a Type-1 hypervisor supporting multiple isolated runtimes. Above Mortise, the Tenon runtime adopts the same modular LibOS principles, enabling architectural consistency and seamless integration. Applications can either run within Tenon or, in some cases, bypass it to interact directly with Mortise for greater efficiency. This LibOS-on-LibOS structure enables flexible, fine-grained composition, rapid adaptation to diverse edge hardware, and improved maintainability and deployment portability.
@@ 5.2 Shared Foundations and a Capability-Based Approach
5.3 LLM-Based Dynamic Adaptation
To support heterogeneous and dynamic edge environments, TenonOS adopts a LLM-based orchestration mechanism that automatically selects and composes micro-libraries at runtime. This approach leverages the semantic reasoning capability of large language models to bridge the gap between high-level objectives and low-level system components, enabling demand-driven, application-specific runtime construction.
The orchestration pipeline proceeds in three major stages:
  1. 1.
  2. 2.
    The resulting representation is a directed acyclic graph where:
  3. 3.
Through this integration of semantic orchestration and graph-based validation, TenonOS can automatically assemble minimal yet sufficient runtime environments tailored to a given workload and hardware context. The graph ensures that only coherent and dependency-consistent libraries are composed, while the LLM provides flexibility in aligning abstract goals with concrete runtime components. This design significantly reduces redundant configurations, prevents invalid or hallucinated options, and enables rapid reconfiguration in edge scenarios.
@@ 5.4 Tenon
LibOS implementations like Graphene [42] and Unikraft [44] have demonstrated their effectiveness in providing lightweight environments for cloud computing. However, embedded systems—especially those in mixed-criticality environments or autonomous driving platforms—also require low-overhead and modular operating systems. Despite these needs, current LibOS designs lack real-time capabilities, which are essential for embedded applications. The absence of deterministic scheduling and reliable timing constraints makes them unsuitable for high-reliability use cases.
Another significant limitation is that most LibOS architectures rely on a single address space (SAS) and support only single-process execution. While this design simplifies implementation and enhances performance for single-application scenarios, it poses challenges when porting multi-process applications. Many general-purpose applications require process isolation, concurrency, and fault separation—features inherent to multi-process models. Adapting them to fit a single-process architecture often demands extensive refactoring, resulting in high migration costs and compatibility issues [45].
@@ 5.5 Mortise
To meet the stringent performance and timing demands of edge workloads, Mortise supports executing applications directly on the virtualization layer, bypassing the guest OS entirely. Applications can be compiled with minimal runtime services and device drivers into a self-contained binary. This image runs directly atop the hypervisor without relying on a guest kernel, enabling fine-grained control over hardware and eliminating OS-level mediation.
@@ 5.5.2 Guest OS Lifecycle Management Library
This dual-mode scheduling design mitigates the inefficiencies of the “double scheduler” problem [46], where both the hypervisor and guest OS independently manage scheduling.
Figure 4: Mode 1 – Static Allocation Bootup: ① The first CPU boots and initializes the Mortsie hypervisor. ② Mortsie reads a static configuration file that specifies the number of instances and their CPU assignments. ③ All remaining CPUs register themselves with Mortsie to participate in the system. ④ Based on the configuration, Mortsie launches Tenon or Linux instances on dedicated CPUs, ensuring strong hardware-level isolation between execution domains.
Mode 1 (Figure 4) enforces static CPU partitioning to ensure temporal and spatial isolation among LibOS instances. At boot, Mortsie reads a configuration file defining the number of instances and their CPU assignments. Each instance is pinned to dedicated physical cores with no sharing, avoiding dynamic scheduling and enabling predictable execution.
@@
To minimize cache contention, Mortsie optionally supports LLC partitioning via page coloring. Though this may increase memory fragmentation and startup latency, it can be enabled per instance based on need.
Figure 5: Mode 2 – Dynamic Allocation Bootup: ① The first CPU boots and initializes the Mortsie hypervisor. ② Mortsie reads system configuration file that defines system policies. ③ All CPUs register with Mortsie, and hypercalls are enabled for dynamic instance control. ④ Based on runtime requests via hypercalls, Mortsie dynamically launches, pauses, or reclaims Tenon instances and assigns CPUs as needed.
@@
5.5.3 OS Isolation and Communication
As multi-OS deployments become more common—particularly in edge environments—efficient inter-VM communication (IVC) becomes essential. Scenarios involving real-time processing or dynamic service composition demand low-overhead channels that support both shared-memory and message-passing patterns. Mortise adopts a LibOS-inspired architecture that flattens the virtualization stack, enabling IVC mechanisms comparable to in-process IPC. By minimizing abstraction between guests and the hypervisor, Mortise supports lightweight, high-throughput communication between OS instances without sacrificing isolation, making it well-suited for collaborative workloads on constrained platforms.
6 Prototype
6.1 Overhead
6.1.1 Code Size and Memory Footprint
We evaluate the code size and memory usage of TenonOS to assess its suitability for resource-constrained deployments. Table 1 summarizes the source lines of code (SLoC) and memory footprint across major components. The entire system comprises only 11,348 lines of code and occupies approximately 361 KB in memory, including all text, data, and BSS sections.
@@ 6.1.1 Code Size and Memory Footprint
This small footprint simplifies verification and validation, minimizes attack surfaces, and enables fast deployment, which are all valuable for embedded or safety-critical settings. Compared to traditional hypervisors such as NOVA (20K LOC) and Bao (16K LOC), our virtualization layer is significantly smaller (13.6K LOC).
This reduction is largely attributed to the LibOS-based architecture, which offloads process and thread management to user-level runtimes. By focusing solely on core virtualization functions—such as resource isolation and hypercall handling—TenonOS avoids the complexity of monolithic hypervisors with built-in schedulers, IPC mechanisms, or device emulation layers.
One of the key benefits of adopting a LibOS-based architecture for the virtualization layer is its significantly reduced code size compared to traditional monolithic hypervisors. By employing a minimalist design, our virtualization layer focuses exclusively on core components, such as resource management and hypercall handling, while delegating functionalities like process and thread management to single-address-space LibOS instances running on top. This modular approach eliminates the need for complex subsystems commonly found in general-purpose hypervisors, such as schedulers and device emulation layers, which can greatly inflate the codebase.
@@ 6.2.1 Baremetal Tenon Scheduling Performance
Table 5: RTOS preempt time
@@
Figure 6 shows a cycle-level comparison of interrupt handling across three RTOSs—Tenon, Zephyr, and RT-Thread—covering three phases: saving interrupt context, dispatching the Interrupt Service Routine (ISR), and restoring context. Tenon demonstrates the lowest total latency at 208 cycles, compared to RT-Thread (335 cycles) and Zephyr (355 cycles), while maintaining balanced performance across all phases. These results indicate that Tenon’s interrupt model, which avoids reliance on generic kernel abstractions, is well-suited for time-sensitive workloads. Its predictable behavior makes it particularly appropriate for embedded systems requiring low and bounded latency.
@@
6.2.2 Baremetal Tenon Real-time Stablity
@@ 6.3 TenonOS Running Efficiency
6.3.1 TenonOS Real-time Effificency
@@ 7 Conclusion
In this work, we proposed TenonOS, a self-generating intelligent software framework that addresses the challenges of modern heterogeneous and dynamic computing environments. By combining lightweight LibOS principles with a modular architecture, TenonOS delivers a flexible and scalable solution capable of adapting to rapidly evolving software demands. At its core, the Mortise hypervisor enhances system adaptability through efficient resource management, strong isolation, and dynamic CPU allocation, supporting both static and dynamic operational modes. Our evaluation demonstrates TenonOS’s superior real-time performance, outperforming traditional RTOS solutions like Zephyr and RT-Thread in scheduling efficiency, latency predictability, and interrupt handling. The micro-library design of Mortise not only reduces the TCB for improved security but also enables high reusability and maintainability with minimal runtime overhead. Furthermore, TenonOS leverages AI-driven decision-making and a unified library pool to dynamically reconfigure itself across diverse hardware and software environments, from resource-constrained embedded systems to high-performance real-time applications. Collectively, these advancements position TenonOS as a next-generation RTOS framework that overcomes the limitations of monolithic architectures, offering a robust foundation for future heterogeneous computing systems.

### v2

@@ 3 Related Work
Operating systems designed for embedded environments must carefully balance flexibility, performance, and strong isolation. Meeting these demands necessitates advancements in both operating system architecture and the underlying virtualization framework. Lightweight, application-specific operating systems—such as Library OSes (LibOS)—minimize overhead and enable precise control over resource usage. Concurrently, minimal type-1 hypervisors incorporating novel architectural designs provide robust isolation and predictable performance, both of which are critical for safety-critical and mixed-criticality applications.
@@ 3.1 Type-1 Hypervisors
Virtualization technology has advanced considerably, with type-1 hypervisors playing a key role in balancing performance, security, and resource isolation. Xen [barham2003xen], a pioneer in type-1 hypervisor design, introduced the classic split-domain architecture. Running directly on bare-metal hardware, Xen separates device management (handled by Dom0) from guest virtual machines (hosted in DomU), ensuring strong isolation and security. Its support for both ARM and x86 architectures enhances compatibility in heterogeneous deployments. Extensions such as RT-Xen further enable latency-sensitive applications in industrial automation and IoT environments. Rust-Shyper [MO2023102948] represents a recent innovation at the programming language level. By leveraging Rust’s memory safety guarantees, Rust-Shyper minimizes vulnerabilities that are common in C/C++ implementations. This approach enhances security and reliability, which is especially important for embedded scenarios such as 5G base stations and distributed AI inference.
Further advancing hypervisor architecture, NOVA [steinberg2010nova] adopts a microhypervisor (microkernel-inspired) strategy, drastically reducing the trusted computing base (TCB) and improving system modularity. By delegating most functionalities to unprivileged user-space components, NOVA achieves strong fault tolerance, making it suitable for embedded systems deployed in untrusted environments like smart city infrastructure and autonomous devices. Continuing this trend towards modularity and determinism, BAO [martins2020bao] is a lightweight, bare-metal hypervisor designed for multi-core embedded systems. BAO applies a modular architecture and statically allocates hardware resources during initialization, achieving deterministic behavior with strong spatial and temporal isolation. This makes BAO particularly suitable for safety-critical and mixed-criticality applications, such as autonomous vehicles and medical devices.
This progression—from Xen’s classic architecture, through Rust-based language innovation, microkernel modularization in NOVA, to BAO’s lightweight and modular design—demonstrates how virtualization technology is evolving towards greater modularity, security, and efficiency for modern embedded computing demands.
Figure 3: The overall architecture of TenonOS. The left section illustrates a traditional monolithic OS-Hypervisor design. The middle section depicts the Librarization process, where system functionalities are modularized into a library pool comprising reusable components. The dynamic orchestration and composition mechanism allows rebuilding and generating TenonOS to adapt to heterogeneous hardware resources, forming a lightweight and agile LibOS on LibOS structure.
@@ 3.2 Flexibility and Performance of Embedded LibOS
LibOS have emerged as a promising paradigm for constructing flexible, lightweight and high-performance software systems. Unlike traditional monolithic operating systems, a LibOS encapsulates only the essential OS abstractions and services required by an application, enabling application-specific optimizations. Early projects such as Exokernel [engler1998exokernel] pioneered the separation of application logic from resource management, allowing applications to directly manage hardware resources.
Recent advancements have focused on integrating LibOS into virtualized environments. For example, Graphene [tsai2017graphene] demonstrated the feasibility of running unmodified applications in isolated LibOS instances. This approach maintains compatibility with existing APIs while enhancing security and performance. In cloud computing, LibOS-based systems such as OSv [kivity2014osv] and Unikraft [kuenzer2021unikraft] have achieved significant improvements in startup latency, memory footprint, and runtime efficiency compared to traditional virtual machines. By leveraging their isolated architecture, these systems minimize overhead and enable further application-specific optimizations.
@@ 4.1 TenonOS Architecture and Design Philosophy
Figure 3 presents the overall architecture of TenonOS, which treats the virtualization layer and guest OS as a unified whole. Traditionally, as shown on the left of the figure, the hypervisor and operating system are organized in a monolithic, tightly coupled stack. In contrast, TenonOS proposes a novel Librarization process, illustrated in the middle of the figure, where both the hypervisor and the OS are dismantled into a unified pool of reusable library components. These modules encapsulate core functionalities, and a dynamic library orchestration mechanism selects and composes the appropriate modules in response to specific hardware and application contexts. The right side of the figure shows the reconstructed modular stack: the bottom layer, Mortise, is a LibOS-based hypervisor that runs directly on hardware and can act as either a traditional OS or dynamically extend to a Type-1 hypervisor supporting multiple isolated runtimes. Above Mortise, the Tenon runtime adopts the same modular LibOS principles, enabling architectural consistency and seamless integration. Applications can either run within Tenon or, in some cases, bypass it to interact directly with Mortise for greater efficiency. This LibOS-on-LibOS structure enables flexible, fine-grained composition, rapid adaptation to diverse hardware, and improved maintainability and deployment portability.
@@ 4.3 Tenon
LibOS implementations like Graphene [tsai2017graphene] and Unikraft [kuenzer2021unikraft] have demonstrated their effectiveness in providing lightweight environments for cloud computing. However, embedded systems—especially those in mixed-criticality environments or autonomous driving platforms—also require low-overhead and modular operating systems. Despite these needs, current LibOS designs lack real-time capabilities, which are essential for embedded applications. The absence of deterministic scheduling and reliable timing constraints makes them unsuitable for high-reliability use cases.
Another significant limitation is that most LibOS architectures rely on a single address space (SAS) and support only single-process execution. While this design simplifies implementation and enhances performance for single-application scenarios, it poses challenges when porting multi-process applications. Many general-purpose applications require process isolation, concurrency, and fault separation—features inherent to multi-process models. Adapting them to fit a single-process architecture often demands extensive refactoring, resulting in high migration costs and compatibility issues [peter2015arrakis].
@@ 4.4 Mortise
To meet the stringent performance and timing demands of time-critical workloads, Mortise supports executing applications directly on the virtualization layer, bypassing the guest OS entirely. Applications can be compiled with minimal runtime services and device drivers into a self-contained binary. This image runs directly atop the hypervisor without relying on a guest kernel, enabling fine-grained control over hardware and eliminating OS-level mediation.
@@ 4.4.2 Guest OS Lifecycle Management Library
This dual-mode scheduling design mitigates the inefficiencies of the “double scheduler” problem [rtxen], where both the hypervisor and guest OS independently manage scheduling.
Figure 6: Mode 1 – Static Allocation Bootup: ① The first CPU boots and initializes the Mortise hypervisor. ② Mortise reads a static configuration file that specifies the number of instances and their CPU assignments. ③ All remaining CPUs register themselves with Mortise to participate in the system. ④ Based on the configuration, Mortise launches Tenon or Linux instances on dedicated CPUs, ensuring strong hardware-level isolation between execution domains.
Mode 1 (Figure 6) enforces static CPU partitioning to ensure temporal and spatial isolation among LibOS instances. At boot, Mortise reads a configuration file defining the number of instances and their CPU assignments. Each instance is pinned to dedicated physical cores with no sharing, avoiding dynamic scheduling and enabling predictable execution.
@@
To minimize cache contention, Mortise optionally supports LLC partitioning via page coloring. Though this may increase memory fragmentation and startup latency, it can be enabled per instance based on need.
Figure 7: Mode 2 – Dynamic Allocation Bootup: ① The first CPU boots and initializes the Mortise hypervisor. ② Mortise reads system configuration file that defines system policies. ③ All CPUs register with Mortise, and hypercalls are enabled for dynamic instance control. ④ Based on runtime requests via hypercalls, Mortise dynamically launches, pauses, or reclaims Tenon instances and assigns CPUs as needed.
@@
4.4.3 OS Isolation and Communication
As multi-OS deployments become more common—particularly in embedded environments—efficient inter-VM communication (IVC) becomes essential. Scenarios involving real-time processing or dynamic service composition demand low-overhead channels that support both shared-memory and message-passing patterns. Mortise adopts a LibOS-inspired architecture that flattens the virtualization stack, enabling IVC mechanisms comparable to in-process IPC. By minimizing abstraction between guests and the hypervisor, Mortise supports lightweight, high-throughput communication between OS instances without sacrificing isolation, making it well-suited for collaborative workloads on constrained platforms.
4.5 LLM-Based Dynamic Adaptation
To support heterogeneous and dynamic embedded environments, TenonOS adopts an LLM-based orchestration mechanism that automatically selects and composes micro-libraries at runtime. This approach leverages the semantic reasoning capability of large language models to bridge the gap between high-level objectives and low-level system components, enabling demand-driven, application-specific runtime construction.
The orchestration pipeline proceeds in three major stages:
  1. 1.
  2. 2.
    • (a)
    • (b)
    The resulting representation is a directed acyclic graph where:
    • (a)
    • (b)
  3. 3.
Through this integration of semantic orchestration and graph-based validation, TenonOS can automatically assemble minimal yet sufficient runtime environments tailored to a given workload and hardware context. The graph ensures that only coherent and dependency-consistent libraries are composed, while the LLM provides flexibility in aligning abstract goals with concrete runtime components. This design significantly reduces redundant configurations, prevents invalid or hallucinated options, and enables rapid reconfiguration in embedded scenarios.
5 Prototype
5.1 Overhead
5.1.1 Code Size and Memory Footprint
@@
We evaluate the code size and memory usage of TenonOS to assess its suitability for resource-constrained deployments. Table 1 summarizes the source lines of code (SLoC) and memory footprint across major components. The entire system comprises only 11,348 lines of code and occupies approximately 361 KiB in memory, including all text, data, and BSS sections.
@@
Compared to traditional hypervisors such as NOVA (20K LOC) and Bao (16K LOC), our virtualization layer is significantly smaller (11.3K LOC). This substantial reduction is largely attributed to the LibOS-based architecture, which delegates process and thread management to single-address-space LibOS instances rather than handling them within the hypervisor. By strictly limiting its scope to core virtualization functions—such as resource isolation and hypercall handling—TenonOS eliminates the need for complex subsystems typically found in monolithic hypervisors, including internal schedulers, IPC mechanisms, and device emulation layers. Consequently, this compact footprint simplifies verification and validation, minimizes the attack surface, and enables fast deployment, all of which are critical for safety-critical embedded environments.
@@ 5.2.1 Baremetal Tenon Scheduling Performance
Table 5: Preemption Times of Different RTOSes
@@
Figure 8 shows a cycle-level comparison of interrupt handling across three RTOSs including Tenon, Zephyr, and RT-Thread, covering three phases: saving interrupt context, dispatching the Interrupt Service Routine (ISR), and restoring context. Tenon demonstrates the lowest total latency at 208 cycles, compared to RT-Thread (335 cycles) and Zephyr (355 cycles), while maintaining balanced performance across all phases. These results indicate that Tenon’s interrupt model, which avoids reliance on generic kernel abstractions, is well-suited for time-sensitive workloads. Its predictable behavior makes it particularly appropriate for embedded systems requiring low and bounded latency.
@@
5.2.2 Baremetal Tenon Real-time Stability
@@ 5.3 TenonOS Running Efficiency
5.3.1 TenonOS Real-time Efficiency
@@
5.4 LLM‑Based Orchestration
Figure 10 compares end‑to‑end generation time and build success for different orchestration strategies on a dataset of 14 Unikraft community–provided application configurations. Pure LLM prompting (DeepSeek‑R1, DeepSeek‑V3) is fast but rarely produces a working Unikraft configuration. Integrated into TenonOS, the orchestrator (TenonOS) achieves the best overall tradeoff, combining high build accuracy with acceptable orchestration latency.
Figure 10: TenonOS generation efficiency and accuracy under different orchestration strategies. The x‑axis shows the end‑to‑end OS generation time, and the y‑axis shows the number of applications successfully built. TenonOS’s orchestration achieves the highest success rate with acceptable orchestration time, outperforming both pure LLM prompting and manual configuration.
@@ 6 Conclusion
In this work, we proposed TenonOS, a self-generating intelligent software framework for time-critical embedded systems that addresses the challenges of modern heterogeneous and dynamic computing environments. By combining lightweight LibOS principles with a modular architecture, TenonOS delivers a flexible and scalable solution capable of adapting to rapidly evolving software demands and diverse criticality levels. At its core, the Mortise hypervisor enhances system adaptability through efficient resource management, strong isolation, and dynamic CPU allocation, supporting both static and dynamic operational modes for co-located real-time and general-purpose workloads. Our evaluation demonstrates TenonOS’s superior real-time performance, outperforming traditional RTOS solutions like Zephyr and RT-Thread in scheduling efficiency, latency predictability, and interrupt handling, while maintaining near bare-metal behavior even in mixed-criticality deployments where Tenon is virtualized by Mortise alongside Linux. The micro-library design of Mortise not only reduces the TCB for improved security but also enables high reusability and maintainability with minimal runtime overhead. Furthermore, TenonOS leverages AI-driven decision-making and a unified library pool to dynamically reconfigure itself across diverse hardware and software environments, from resource-constrained embedded SoCs to high-performance real-time applications. Collectively, these advancements position TenonOS as a next-generation RTOS framework that overcomes the limitations of monolithic architectures, offering a robust foundation for future heterogeneous and mixed-criticality computing systems.